An indoor heater operating on the Carnot cycle is warming the house up at a rate of 30 kJ/s to maintain the indoor temperature at 72 ºF. What is the power operating the heater if the outdoor temperature is 30 ºF?

Physical Chemistry Question!

Please provide complete calculations. Answer will be rated!
Answer #1

T2 = 30°F = 272.039K T1 = 295.372K

Efficiency of heater = 1- T2/T1 = 1- (272.039/295.372) = 0.0789953

The power operating the heater = 30kJ/s ÷ 0.0789953

= 379.77 kJ/s

= 380 kJ/s